How many grams of ammonia chloride must be added to 0.05kg of water to produce a solution that freezes at -5°C? The cryoscopic constant of water is 1.86°C kg mol-1. The molar mass of ammonium chloride is 53.5gmol-1
1
Expert's answer
2020-06-08T15:49:49-0400

WB = unknown

MB = 53.5 g/mol

WA = 0.05 kg

∆T = 5 °C

Kf = 1.86 °C×kg/mol

∆T = (Kf×WB)/(MB×WA)

WB = (∆T×MB×WA)/Kf

WB = (5×53.5×0.05)/1.86 = 7.19 g

Answer: 7.2 g